Ethylene glycol C2H6, is a common automobile antifreeze agent. Calculate the freezing point of solution containing 651 grams of EG in 2505 grams of water. The molar mass of EG is 62.01 g/mol (Kf=1.86° C/m) and (Kb= 0.52 ° C/m for water)


Show your solution.

1
Expert's answer
2022-02-24T00:40:02-0500

Depression in any solute is known as the colligative property.


Here, ΔTf = kf*m m = molality which is equal to moles of solute per kg solvent


Moles of ethylene glycol = 12.4/62


Molality = 0.2/0.1=2


By substituting we get,


ΔTf = 1.86*2 = 3.72K


Difference in freeezing point = 3.72K ie Depression


Boiling point elevation = ΔTb = Kb*b*i


=0.512*2 = 1.024K


Boiling point of solution = 373+1.024= 374.024k= 101.024deg C


The boiling point increases with increase in ethylene glycol percentage.


The use of ethylene glycol increases boiling point and depresses the freezing point.


It act as a anti freezing agent also and can be keep this in car radiator during summer